A is 5 years older than B. The product of their ages is 84. How old is each
Answer by farbonara(13)

A-B=5 ---(1)
AxB=84 ---(2)
Rearranging (1) making A the subject,
A=B+5 ---(3)
Substitute (3) into (2),
(B+5)xB=84
B^2+5B=84
B^2+5B-84=0
(B-7)(B+12)=0
B=7 or B=-12(reject)
Using (3),
A=7+5=12